If 7.3 mL of 1.25 M 𝐻𝑁𝑂_3 is required to neutralize 25.00 mL of a potassium hydroxide solution, what is the molarity of the potassium hydroxide?


𝐻𝑁𝑂_3 + KOH →KN𝑂_3 + 𝐻_2 𝑂


1
Expert's answer
2022-01-26T02:27:01-0500

Mole ratio of


HNO3 ; KOH"=1:1"


"\frac{7.3}{1000}×1.25=\frac{25}{1000}×M_k"


"M_k=\frac{7.3×1.25}{25}"


"0.365M"